OpenStudy (anonymous):

In less modern times, door-to-door salesmen carried samples of large-sized appliances, such as stoves and washers. Now an antique, one such sample is 14 inches tall. The scale of the sample is . What was the actual height of the appliance?

OpenStudy (anonymous):

Scale sample: 1 inch: 2/7 foot

OpenStudy (anonymous):

does anyone have any idea how to do this???

OpenStudy (anonymous):

a. 49 b.4 c.2 d. 48

OpenStudy (anonymous):

I only have 8 more questions on my hw plz help

OpenStudy (anonymous):

use a proportion 1/ 2/7 = x/14 14 = 2/7x multiply each side by the reciprocal of 2/7 which is 7/2 and you get 49... A

OpenStudy (anonymous):

that wrong

OpenStudy (anonymous):

oh, my bad its B...i did it wrong

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

need help on how to do it

OpenStudy (anonymous):

use a proportion 1 / 2/7 = 14/x x = 14 * 2/7 x = 4

OpenStudy (anonymous):

maybe easier to try \[\frac{2}{7}=\frac{x}{14}\] \[x=\frac{2\times 14}{7}=2\times 2=4\] keep the variable in the numerator

OpenStudy (anonymous):

or put more simply, each inch represents \[\frac{2}{7}\] of a foot, so 14 inches represents \[14\times \frac{2}{7}\] of a foot